OpenStudy (anonymous):

Im Confused !! The height of an equilateral triangular prism decreases by two units. The new volume is less than the original by how much? a.two times the area of the base b.two times the length of the height c.two more than the the area of the base d.two more than the length of the height

OpenStudy (cwrw238):

vol = area of base * height = Ah new vol = A(h - 2) = Ah - 2A

OpenStudy (cwrw238):

2A = two times area of base

OpenStudy (cwrw238):

ok?

OpenStudy (anonymous):

OK .